# HA-RDet: Hybrid Anchor Rotation Detector for Oriented Object Detection

### v1

@@ 1 Introduction
Existing oriented detectors [5, 26, 9, 25, 3] typically rely on proposal-driven frameworks utilizing a Region Proposal Network (RPN), which employs numerous anchors with varying scales and aspect ratios to generate Regions of Interest (RoIs). However, horizontal RoIs often cover multiple instances, causing ambiguity, while rotated RoIs improve recall but incur high computational costs. Hence, developing an efficient region proposal network capable of generating high-quality proposals is crucial to overcome accuracy and computational limitations in current state-of-the-art oriented detectors. We identify two key reasons for the lack of efficiency in region proposal-based oriented detectors: (1) Most oriented object detection methods adopt Anchor-based or Anchor-free training approaches, each with its own challenges and trade-offs in terms of performance and computational efficiency. Anchor-based methods achieve good performance but are sensitive to oriented anchor hyperparameters, while Anchor-free methods accelerate training speed but suffer from decreased detection performance due to a shortage of level features. (2) The conventional convolution technique, designed with fixed receptive fields and aligned with axes, is ill-suited for objects with arbitrary orientations and varying scales in aerial images. Predefined anchor representations often fail to fully capture the shape and orientation of objects, leading to misalignment between fixed convolution features and object orientations, thus hindering effective object localization and classification in oriented object detection models.
@@
This study introduces a novel method called Hybrid-Anchor Rotation Detector (HA-RDet) for detecting oriented objects, which combines the strengths of Anchor-based and Anchor-free mechanisms through a Hybrid anchor training approach, employing a single anchor per location on the feature map. In the anchor-based component, we introduce a new Orientation-aware Convolution (O-AwareConv) technique that enhances proposal quality by adapting to object shape and orientation. Through extensive experimentation on established oriented object detection datasets and comparison with previous detectors, we demonstrate the effectiveness of our approach. Our method addresses challenges such as diminishing training samples and over-encoding information by reducing anchor hyperparameters initially and refining them later to generate high-quality proposals. Additionally, we design a lightweight transformation network in the final R-CNN head to extract rotated proposals from horizontal ones. Unlike prior studies, HA-RDet exhibits superior efficiency in both training and testing, requiring fewer computational resources while achieving promising accuracy and speeding up detection. This work aims to offer an elegant and efficient region proposal-based oriented detector that bridges the gap between anchor-free and anchor-based approaches, serving as a robust baseline for future investigations.
@@
  • Designing a simple yet effective light-weight network for orientation transformation
@@ 3 Proposed Hybrid-Anchor Rotation Detector
Most oriented object detectors follow the anchor-based strategy for predicting the heavily encoded oriented bounding boxes. They often generate multiple anchors of different ratios and scales to allocate many instances of various shapes and sizes greedily. This approach requires a large number of anchors, leading to high computational requirements. On the other hand, the anchor-free mechanism uses only a single anchor per location on a feature map, but it suffers from accuracy deterioration due to the lack of positive training samples. Many studies improve the performance of either method via post-refining, new assigner strategies, new anchor definitions, etc. Two strategies standstill and need to be fully exploited in oriented object detection.
@@ 3.1 Hybrid-Anchor RPN
Anchor encoder-decoder. We first investigate HA-RDet adopting RetinaNet [14], initially designed for generic horizontal object detection but adapted for oriented object detection, where anchors are encoded as 5D $(x,y,w,h,\theta)$. Utilizing the Anchor-free assigner, the Anchor-free head aims to train the oriented anchor generator from feature maps extracted via regular 2D convolution (2DConv). Subsequently, the Anchor-based head employs a 5D encoder-decoder strategy to regress offset values $(\delta x,\delta y,\delta w,\delta h,\delta\theta)$ for the final oriented proposals. However, this experimental approach proves ineffective due to two primary reasons: (1) Insufficient positive samples hinder the model from acquiring adequate information, resulting in insufficient oriented proposals; (2) Excessive information encoded within the oriented anchors in the Anchor-free head prevents the anchor generator from converging.
@@
where $\delta=(\delta_{x},\delta_{y},\delta_{w},\delta_{h})$ and $\widehat{\delta}=(\widehat{\delta_{x}},\widehat{\delta_{y}},\widehat{\delta_{w}},\widehat{\delta_{h}})$ are predicted encoded value and target encoded value, respectively. Having regressed $\widehat{\delta}$, we can easily decode the proposal $(a^{\prime}_{x},a^{\prime}_{y},a^{\prime}_{w},a^{\prime}_{h})$. Having the classification branch omitted, the Anchor-free head aims to minimize the bounding box loss function $L_{AF}$:
@@
where $w^{AF}$ is the loss weight used for fine-tuning the model.
Anchor-Based head. The Anchor-based head, defined by Eq. 1, receives anchors generated from the preceding stage along with their extracted features from Orientation-aware Convolution. Its objective is to generate refined, high-quality proposals. The Anchor-based bounding box loss function $L_{AB}$ is derived from Eq. 2, albeit with a different weight $w^{AB}$. However, different from Anchor-free, the classification branch with binary cross-entropy loss in this stage is used. The overall loss of our Region Proposal Network is the sum of the Anchor-free loss $L_{AF}$ and the Anchor-based loss $L_{AB}$, denoted as $L_{RPN}$.
$L_{RPN}=L_{AF}+L_{AB}$ (3)
Orientation-aware Convolution. Object detection in aerial images presents unique challenges compared to standard perspective images due to densely populated instances, diverse aspects, and varying scales. Models tailored for aerial object detection require precise localization and robust feature extraction.
Our introduced Orientation-Aware Convolution (O-AwareConv) is utilized in anchor-based heads, aiming to achieve global feature alignment by adaptively considering the object’s ground truth orientation $\theta_{GT}$. The proposed method involves designing a convolution that extensively calculates the shape offset value for each horizontal anchor in conjunction with the corresponding ground-truth orientation offset value. During training, a multi-level horizontal anchor list is generated, and O-AwareConv computes the shape offset field $(x,y,w,h)$ for each anchor to adjust the aligned features’ sampling locations. Additionally, orientation offset values $R(\theta_{GT})^{T}$ are adapted based on the corresponding ground-truth orientation $\theta_{GT}$, where $T$ denotes matrix transposition. This process allows convolutional features at a given location to be transformed into oriented features based on the specific anchor and ground-truth target, as depicted in Fig. 2. In summary, for each location $p$ on a feature map $X$, O-AwareConv considers the corresponding anchor $a^{*}=(x,y,w,h)$, stride $S$, kernel weight $w$, and kernel size $k$, which is represented as Eq. 4:
@@
In our approach, each anchor corresponds positively to only one specific ground truth, whereas each ground truth may be associated with multiple anchors. Leveraging every training sample, or training anchor, our method retains negative anchors in anchor-free heads as a set of potential predefined candidates, with the expectation of gradual improvement during the refining stage. During testing, where corresponding ground truth is unavailable, the orientation offset field is set to zero when calculating the final features offset.
@@ 1. A = anchor_generator(X)
 2. target = assigner(A, rgt)[’index’]
@@ 5. A = decode(A, bbox_pred)
 7. target = assigner(A, rgt)[’index’]
@@ 15. A = decode(A, bbox_pred)
17. loss_cls = class_loss(bbox_cls, label[target])
@@
Algorithm 1 Hybrid-Anchor RPN
Compared to regular 2D convolution, which assumes objects in an image align with the axes, our approach utilizes Deformable Convolution (DeformConv) [2] with augmented offsets. However, DeformConv may inaccurately sample from incorrect locations, particularly with weak supervision for densely distributed objects. To address this limitation, we introduce the Orientation-Aware Convolution during the training phase for every generated horizontal anchor, ensuring global convergence of the Hybrid-Anchor RPN. Unlike DeformConv and 2DConv, offset field in O-AwareConv is inferred from the actual shape and orientation of objects. Fig. 2 demonstrates that our O-AwareConv accurately extracts features. In inference, global features are extracted effectively, enhancing anchor prediction accuracy without the need for the orientation feature channel.
@@ 4.1 Datasets
DOTA-v1.0 [24] is a large-scale dataset comprising high-quality aerial images for detecting objects, containing $2,806$ high-defined images with resolutions $4000\times 4000$. It classifies annotated instances into $15$ common categories: plane (PL), baseball diamond (BD), bridge (BR), ground track field (GTF), small vehicle (SV), large vehicle (LV), ship (SH), tennis court (TC), basketball court (BC), storage tank (ST), soccer ball field (SBF), roundabout (RA), harbor (HA), swimming pool (SP) and helicopter (HC). During the training and testing process, the images are divided into cropped patches of size $1024\times 1024$ pixels with a stride of $824$ pixels, including random flipping is applied as a form of data augmentation.
@@
Table 1: Comparison of SOTA results on the oriented object detection DOTA 1.0 task.
@@
Table 4: Models trained on DOTA with an anchor’s scale larger than 6 perform better on large-instance classes such as GTF, HA, HC, PL, etc. However, small-instance classes, such as SV, LV, SH, etc., have higher performance on models with the anchor’s scale of less than 6.
@@ 4.3 Comparisons with State-of-the-Arts
The experimental results presented in Tab. 1 demonstrate the performance of HA-RDet compared to previous detectors on the DOTA-v1 dataset. HA-RDet achieves state-of-the-art asymptotic results with 75.41 mAP and 76.02 mAP using the regular ResNet50 and ResNet101 backbones, respectively. The highest result of 77.01 mAP was particularly impressive with ResNeXt101_DCNv2, surpassing most existing one-stage and anchor-free methods. By utilizing O-Aware convolution and aligning features accordingly, our model significantly improves the accuracy of detecting irregularly shaped objects such as bridges, storage tanks, and harbors, efficiently addressing high aspect ratio challenges.
@@
Tab. 3 presents the outcomes obtained from the DIOR-R dataset, where our novel Hybrid-Anchor training schemes approach demonstrates superior performance. Using the backbone ResNet50, our method achieves the highest accuracy at 65.3 mAP, surpassing the well-known two-stage Oriented R-CNN by 0.2 mAP, the one-stage $S^{2}$-Net by 3.6 mAP, and the Anchor-free Oriented Proposal Generator (AOPG) by 0.9 mAP.
We conduct further experiments to compare our proposed Hybrid-Anchor training scheme with the well-known anchor-free $S^{2}$-Net and anchor-based Oriented R-CNN methods in Tab. 5. We evaluate these methods beyond the mean Average Precision (mAP) measure, such as inference speed (FPS), computational resources (VRAM) for storing as well as processing the intermediate data during the training process, and the number of model parameters. The $S^{2}$-Net reached the detection result of 74.19 mAP with a fast inference speed of 15.5 FPS; it consumed 4.6 GB of VRAM due to its relatively low number of parameters. Meanwhile, Oriented R-CNN, with 20 anchors in the training stage, attained a slightly better accuracy than $S^{2}$-Net but with a trade-off regarding inference speed (13.5 FPS) and higher computational resource requirements (14.2 GB of VRAM). Our HA-RDet follows the Hybrid-Anchor training approach and generates only one anchor per location, similar to $S^{2}$-Net. It achieved better detection accuracy (75.41 mAP) while maintaining a stable inference speed of 12.1 FPS. Compared to Oriented R-CNN, our model has a higher parameter requirement for processing but consumes approximately half the amount of VRAM (6.8 GB) due to the significant reduction in the number of generated anchors. Oriented R-CNN utilized dense generated anchors and a complex encoding strategy; therefore, its accuracy and inference speed is slightly higher than our model, but it comes at the cost of significantly higher computational resource requirements. Some visualization of HA-RDet are demonstrated in Fig. 3
@@
Table 5: Evaluation results of HA-RDet with $S^{2}$-Net and Oriented R-CNN on the ResNet50+FPN on DOTA
@@ 4.4 Ablation studies
Effectiveness of O-AwareConv. As mentioned in section 3.1, we introduced a groundbreaking approach called O-AwareConv that enhances the awareness of ground-truth orientation. Tab. 6 compares O-AwareConv against alternative methods under the same configurations to demonstrate its effectiveness in oriented object detection. Our HA-RDet employs Standard Convolution and reaches only 74.1 mAP because the fixed receptive field fails to extract features adequately. On the other hand, Deformable Convolution addresses feature misalignment, yet its performance degrades when detecting densely packed and arbitrarily oriented objects in aerial images. O-AwareConv significantly enhances the performance of our proposed detector, surpassing Deformable Convolution by 0.91 mAP and outperforming standard Convolution by 1.31 mAP.
@@
Table 7: Experimental evaluation of our Hybrid-Anchor RPN by taking out a couple of modules, implemented on ResNet50+FPN on DOTA-v1.
@@
Study on anchor scale. We use one preset anchor with one scale and ratio for each location on a feature map following the anchor-free mechanism Tab. 4. Then, we extend the training longer, and the diminishing of samples happens in models having anchor scale larger than 6 due to the overwhelming number of small-sized instances of the datasets. Finally, we decided to choose the anchor scale of 4 with ratio of 1.0 as the representative for our HA-RDet baseline because of its stability, easy convergence, and suitability for high-quality aerial images.
5 Conclusions
We propose the Hybrid-Anchor training scheme detector called HA-RDet with the novel O-AwareConv to tackle the issues of limited positive samples and over-encoding information in oriented object detection. We extensively experiment on three challenging aerial image datasets, achieving competitive accuracy and inference time against state-of-the-art methods with negligible computational resources.

### v2

@@ 1 Introduction
Existing oriented detectors [5, 30, 9, 29, 3] typically rely on proposal-driven frameworks utilizing a Region Proposal Network (RPN), which employs numerous anchors with varying scales and aspect ratios to generate Regions of Interest (RoIs). However, horizontal RoIs often cover multiple instances, causing ambiguity, while rotated RoIs improve recall but incur high computational costs. Hence, developing an efficient region proposal network capable of generating high-quality proposals is crucial to overcome accuracy and computational limitations in current state-of-the-art oriented detectors. We identify two key reasons for the lack of efficiency in region proposal-based oriented detectors: (1) Most oriented object detection methods adopt Anchor-based or Anchor-free training approaches, each with its own challenges and trade-offs in terms of performance and computational efficiency. Anchor-based methods achieve good performance but are sensitive to oriented anchor hyperparameters, while Anchor-free methods accelerate training and inference speed but suffer from decreased detection performance due to shortage number of preset training anchors. (2) The conventional convolution technique, designed with fixed receptive fields and aligned with axes, is ill-suited for objects with arbitrary orientations and varying scales in aerial images. Predefined anchor representations often fail to fully capture the shape and orientation of objects, leading to misalignment between fixed convolution features and object orientations, thus hindering effective object localization and classification in oriented object detection models.
@@
This study presents a novel method, Hybrid-Anchor Rotation Detector (HA-RDet), for oriented object detection. HA-RDet combines the strengths of anchor-based and anchor-free paradigms through a hybrid anchor training scheme that employs a single anchor per location on the feature map. At the core of our design is the proposed Orientation-Aware Convolution (O-AwareConv), which serves as a bridge between one-stage and two-stage detection heads. O-AwareConv improves proposal quality by dynamically adapting to the object’s shape and orientation. Extensive experiments on standard oriented object detection benchmarks demonstrate the effectiveness of our approach. HA-RDet addresses key challenges such as insufficient positive training samples and over-encoding of geometric information by initially reducing anchor hyperparameters and subsequently refining proposals in later stages. Unlike prior methods, HA-RDet achieves a superior balance between accuracy and computational efficiency in both training and inference. It requires fewer resources while maintaining competitive performance, thereby accelerating the detection process. Overall, HA-RDet offers an elegant and efficient region proposal-based solution that bridges the gap between anchor-based and anchor-free approaches, providing a strong baseline for future research in oriented object detection.
@@ 3 Proposed Hybrid-Anchor Rotation Detector
Most oriented object detectors follow the anchor-based strategy for predicting the 5-D encoded oriented bounding boxes. They often generate multiple anchors of different ratios and scales to allocate many instances of various shapes and sizes greedily. This approach requires a large number of anchors, leading to high computational requirements. On the other hand, the anchor-free mechanism uses only a single anchor per location on a feature map, but it suffers from accuracy deterioration due to the lack of positive training samples. Many studies improve the performance of either method via post-refining, new assigner strategies, new anchor definitions, etc. Two strategies standstill and need to be fully exploited in oriented object detection.
@@ 3.1 Hybrid-Anchor RPN
Anchor encoder-decoder strategy. We begin by examining the vanilla baseline of HA-RDet, which builds on RetinaNet [16]—a detector originally developed for generic horizontal object detection—adapted here for oriented object detection by representing anchors in a 5D format $(x,y,w,h,\theta)$. The Anchor-free head, guided by an Anchor-free assigner, is trained to generate oriented region proposals directly from feature maps produced by standard 2D convolution (2DConv). In the next stage, the Anchor-based head refines these anchors using a 5D encoder-decoder to regress offsets $(\delta x,\delta y,\delta w,\delta h,\delta\theta)$, resulting in the final oriented object proposals. However, this experimental setup proves ineffective for two main reasons, both of which compromise proposal quality and consequently degrade the performance of the anchor-based head: (1) Limited positive training samples results in insufficient and unreliable proposals from the anchor-free generator; (2) Overly complex five-dimensional anchor encoding in the anchor-free head introduces noise and reduces proposal precision.
@@
where $\delta=(\delta_{x},\delta_{y},\delta_{w},\delta_{h})$ and $\widehat{\delta}=(\widehat{\delta_{x}},\widehat{\delta_{y}},\widehat{\delta_{w}},\widehat{\delta_{h}})$ are predicted encoded value and target encoded value, respectively. Having regressed $\widehat{\delta}$, we can easily decode the proposal $(a^{\prime}_{x},a^{\prime}_{y},a^{\prime}_{w},a^{\prime}_{h})$. With the classification branch omitted, the Anchor-free head aims to minimize the bounding box loss function $L_{AF}$:
@@
where $w^{AF}$ is the loss weight used to fine-tune the model. At the end of this stage, from a limited number of preset anchors, we obtain a preliminary set of horizontal candidates, prepared for further refinement.
Orientation-aware Convolution. To extract meaningful geometric cues from the set of predicted candidates, we propose orientation-aware convolution (O-AwareConv), which improves the alignment of global features and allows the extraction of orientation-sensitive features for horizontal anchor candidates. O-AwareConv adaptively incorporates the orientation of the ground truth object $\theta_{GT}$ during training. Specifically, it calculates a shape offset field for each horizontal anchor, in combination with an orientation offset derived from the rotation matrix $R(\theta_{GT})^{T}$, where $T$ denotes the transposition of the matrix. These offsets guide the convolution to adjust its sampling locations, enabling it to align with the true orientation and geometry of objects. As illustrated in Fig. 2, this design transforms conventional convolutional features into orientation-aware representations, tailored to each anchor-ground-truth pair.
Formally, for each spatial location $p$ on a feature map $X$, O-AwareConv uses the anchor candidate $a^{*}=(x,y,w,h)$, stride $S$, kernel weights $W$, and kernel size $k$ to calculate the position $\Omega$ of the final compute feature $Y$ as defined in Eq. 3:
@@
In our approach, each anchor is assigned to at most one ground-truth object, while each ground truth can correspond to multiple anchors. To make full use of the training data, we retain all anchor candidates, including negative ones, in the previous anchor-free heads as a set of potential predefined candidates, anticipating that their quality will improve progressively during the refinement stage. During inference, when ground-truth annotations are not available, the orientation offset field is set to zero, ensuring deterministic feature offset computation.
@@ 1. A = anchor_generator(X)
 2. target = assigner(A, rgt)[‘index’]
@@ 5. A = decode(A, bbox_pred)
 7. target = assigner(A, rgt)[‘index’]
@@ 14. A = decode(A, bbox_pred)
16. loss_cls = objectness(bbox_cls, label[target])
@@
Algorithm 1 Hybrid-Anchor training strategy
Compared to standard 2D convolution, which assumes objects are axis-aligned, our method incorporates Deformable Convolution (DeformConv) [2] with augmented offsets to improve spatial adaptability. However, DeformConv can still sample from suboptimal locations, especially under weak supervision and in densely populated scenes. To mitigate this, our O-AwareConv is used for feature extraction during training for each generated horizontal anchor, enabling robust orientation-aware alignment within the Hybrid-Anchor RPN. Unlike 2DConv and DeformConv, the offset field in O-AwareConv is explicitly derived from the object’s shape and orientation. As shown in Fig. 2, this mechanism allows for more accurate and geometry-aligned feature extraction. During inference, O-AwareConv effectively captures global features without relying on orientation supervision, thereby improving anchor prediction accuracy. At the end of this stage, we obtain a set of horizontal candidates from the Anchor-free head, along with their corresponding orientation-aware features $Y$ extracted by O-AwareConv.
Anchor-Based head. The anchor-based head serves as a refinement module. It re-encodes the input according to Eq. 1, taking as input the anchors generated in the previous stage along with their features extracted by the orientation-aware convolution. Its goal is to produce refined, high-quality proposals. The Anchor-based bounding box loss function $L_{AB}$ is derived from Eq. 2, albeit with a different weight $w^{AB}$. However, unlike Anchor-free, the classification branch with binary cross-entropy loss is used in this stage to estimate the objectness of the predicted region proposal. The overall loss of our Region Proposal Network is the sum of the Anchor-free loss $L_{AF}$ and the Anchor-based loss $L_{AB}$, denoted as $L_{RPN}=L_{AF}+L_{AB}$. Subsequently, we obtain the final set of horizontal region proposals, which effectively capture orientation-sensitive objects and are ready to be transformed into oriented proposals.
@@ 4.1 Datasets
DOTA-v1.0 [28] is a large-scale dataset comprising high-quality aerial images for detecting objects, containing $2,806$ high-defined images with resolutions $4000\times 4000$. It classifies annotated instances into $15$ common categories: plane (PL), baseball diamond (BD), bridge (BR), ground track field (GTF), small vehicle (SV), large vehicle (LV), ship (SH), tennis court (TC), basketball court (BC), storage tank (ST), soccer ball field (SBF), roundabout (RA), harbor (HA), swimming pool (SP) and helicopter (HC). During the training and testing process, the images are divided into cropped patches of size $1024\times 1024$ pixels with a stride of $824$ pixels, including random flipping is applied as a form of data augmentation. We submit our test set results to the evaluation website.
@@
Table 1: Comparison of SOTA results on the oriented object detection DOTA 1.0 task. We report the results of methods using the same training configuration on DOTA 1.0
@@
Table 4: Models with anchor scales above 6 perform better on large objects, while smaller scales yield higher accuracy for small instances
@@ 4.3 Comparisons with State-of-the-Arts
The experimental results in Tab. 1 highlight the performance of HA-RDet on the DOTA-v1 dataset in comparison to prior detectors. While recent methods leveraging ViT-based backbones [25, 15, 33], end-to-end training paradigms [35] and different encoding-decoding [34] mechanisms have achieved superior performance—surpassing both traditional one-stage and two-stage detectors as well as our HA-RDet—this study focuses on fair comparisons under the same configuration (Oriented-RCNN, RoI Transformer, S2ANet). Specifically, we evaluate HA-RDet against methods using the same backbone architecture and data augmentation strategies. Future work on these adjustments would be considered.
Under these settings, HA-RDet demonstrates prominent performance, achieving 75.41 mAP and 76.02 mAP with ResNet-50 and ResNet-101 backbones, respectively. The highest accuracy of 77.01 mAP is attained using ResNeXt101-DCNv2. HA-RDet surpasses most anchor-free methods and delivers results competitive with the well-established two-stage Oriented R-CNN. The incorporation of Orientation-Aware Convolution enables HA-RDet to better align features with object geometry, leading to significantly improved detection accuracy for irregularly shaped objects—such as bridges, storage tanks, and harbors—effectively addressing the challenges posed by high aspect ratios.
@@
Tab. 3 presents the results obtained from the DIOR-R dataset, where our novel Hybrid-Anchor training schemes approach demonstrates superior performance. Using the backbone ResNet50, our method achieves the highest accuracy at 65.3 mAP, surpassing the well-known two-stage Oriented R-CNN by 0.2 mAP, the one-stage $S^{2}$-Net by 3.6 mAP, and the Anchor-free Oriented Proposal Generator (AOPG) by 0.9 mAP.
We conduct further experiments to compare our proposed Hybrid-Anchor training scheme with the well-known anchor-free $S^{2}$-Net and anchor-based Oriented R-CNN methods in Tab. 5. We evaluate these methods beyond the mean Average Precision (mAP) measure, such as inference speed (FPS), computational resources (VRAM) for storing as well as processing the intermediate data during the training process, and the number of model parameters. The $S^{2}$-Net reached the detection result of 74.19 mAP with a fast inference speed of 15.5 FPS; it consumed 4.6 GB of VRAM due to its relatively low number of parameters. Meanwhile, Oriented R-CNN, with 20 anchors in the training stage, attained a slightly better accuracy than $S^{2}$-Net but with a trade-off regarding inference speed (13.5 FPS) and higher computational resource requirements (14.2 GB of VRAM). Our HA-RDet follows the Hybrid-Anchor training approach and generates only one anchor per location, similar to $S^{2}$-Net. It achieved better detection accuracy (75.41 mAP) while maintaining a stable inference speed of 14.9 FPS. Compared to Oriented R-CNN, our model has a higher parameter requirement for processing but consumes approximately half the amount of VRAM (6.9 GB) due to the significant reduction in the number of generated anchors. Oriented R-CNN utilized dense generated anchors and a complex encoding strategy; therefore, its accuracy and inference speed is slightly higher than our model, but it comes at the cost of significantly higher computational resource requirements. Some visualization of HA-RDet are demonstrated in Fig. 3
@@
Table 5: Evaluation results of HA-RDet with $S^{2}$-Net and Oriented R-CNN on the ResNet50+FPN on DOTA. Anchors are saved under CUDA tensors for fast computation
@@ 4.4 Ablation studies
Effectiveness of O-AwareConv. As mentioned in section 3.1, we introduced an approach called O-AwareConv that enhances the awareness of ground-truth orientation. Tab. 6 compares O-AwareConv against alternative methods under the same configurations to demonstrate its effectiveness in oriented object detection. Our HA-RDet employs Standard Convolution and reaches only 74.1 mAP because the fixed receptive field fails to extract orientation-sensitive features. On the other hand, Deformable Convolution addresses feature misalignment, yet its performance degrades when detecting densely packed and arbitrarily oriented objects in aerial images. O-AwareConv significantly enhances the performance of our proposed detector, surpassing Deformable Convolution by 0.91 mAP and outperforming standard Convolution by 1.31 mAP.
@@
Table 7: Experimental evaluation of our Hybrid-Anchor RPN by taking out a couple of modules, using ResNet50+FPN on DOTA.
@@
Study on anchor scale. We use one preset anchor with one scale and ratio for each location on a feature map following the anchor-free mechanism Tab. 4. Then, we extend the training longer, and the diminishing of samples happens in models having anchor scale larger than 6 due to the overwhelming number of small-sized instances of the datasets. Finally, we decided to choose the anchor scale of 4 with ratio of 1.0 as the representative for our HA-RDet baseline given its stability, easy convergence, and suitability.
5 Discussion
We propose HA-RDet, a Hybrid-Anchor training scheme that integrates the novel O-AwareConv to tackle two critical challenges in oriented object detection: the scarcity of positive samples and the risk of over-encoding geometric information. HA-RDet bridges the gap between one-stage and two-stage detection paradigms, striking a balance between training/inference efficiency and detection accuracy. Our method is particularly well-suited for Earth Observation applications, where aerial and satellite imagery often contain densely packed, arbitrarily oriented objects. Use cases include maritime monitoring, urban development assessment, and disaster response. The ability to detect rotated objects precisely is crucial in these contexts, where misalignment can lead to misclassification or missed detections. HA-RDet’s computational efficiency, achieved via a reduced anchor set and lightweight orientation-aware refinement, makes it deployable on edge devices such as drones or low-power remote sensing platforms. Notably, HA-RDet has been quantized and deployed on campus drone systems, enabling continuous monitoring and autonomous governance.